Interest on Bond Investments

On February 1, Hansen Company purchased $66,000 of 8%, 15-year Knight Company bonds at their face amount plus one month’s accrued interest. The bonds pay interest on January 1 and July 1. On October 1, Hansen Company sold $24,000 of the Knight Company bonds acquired on February 1, plus three months’ accrued interest. On December 31, three months’ interest was accrued for the remaining bonds.

Determine the interest earned by Hansen Company on Knight Company bonds for the year. Do not round interim calculations. If required, round final answer to the nearest dollar.

_____
Answer #1

Answer:

Cost of bonds = $66000, rate of interest = 8%

Interest received on july 1 for six months

= 66000 x 8% x 6/12 = $2640

But net income earned is for 5 month as 1 month is accrued at the time of purchase

Therefore, interest earned upto July 1 = 2640 x 5/6 = 2200

Cost of bonds sold on October 1 is $24000

Interest earned on these bonds for 3 months

= 24000 x 8% x 3/12 = $480

Interest earned on remaining bonds for 6 months

= 42000 x 8% x 6/12 = $1680

Total interest earned in the year = 2200+ 480+ 1680

= $4360